OpenStudy (anonymous):

Solve the given equation and indicate the number of values of θ sach that 0≤θ<2π ,that satisfy the iquation; Sina=½ 3^½

OpenStudy (amistre64):

Is the equation: sin(a) = sqrt(3)/2 ? There are 2 angles that have a positive (sin) of sqrt(3)/2 in the unit circle. To solve the equation: we use the (sin inverse) functions; they take a number and spit back an angle. sin(a) = sqrt(3)/2 sin-1(sin(a)) = sin-1(sqrt(3)/2) a = sin-1(sqrt(3)/2) a = pi/3 Now the issue here is that sin-1 only gives you an angle between (-pi/2) and (pi/2) so you have to revert back to the original question to ALL the angles. 60 degrees (pi/3) and, 120 degrees (2pi/3) Does that make sense?

OpenStudy (anonymous):

Equation is sina=plus,minus 1/2 sqrt(3)?

OpenStudy (amistre64):

If it is sin(a) = +- sqrt(3)/2 then there are 4 angles that we can use; and they all have a reference angle of (60 degrees). In quadrant 1 we have (pi/3) which is 60 degrees In quadrant 2 we have (2pi/3) which is 120 degrees In quadrant 3 we have (4pi/3) which is 240 degrees In quadrant 4 we have (5pi/3) which is 320 degrees

OpenStudy (anonymous):

Make sense.Thank you